# SARDP_supp

### v1

@@ Abstract
Spatial autoregressive (SAR) models are important tools for studying network effects. However, with an increasing emphasis on data privacy, data providers often implement privacy protection measures that make classical SAR models inapplicable. In this study, we introduce a privacy-protected SAR model with noise-added response and covariates to meet privacy-protection requirements. However, in this scenario, the traditional quasi-maximum likelihood estimator becomes infeasible because the likelihood function cannot be formulated. To address this issue, we first consider an explicit expression for the likelihood function with only noise-added responses. However, the derivatives are biased owing to the noise in the covariates. Therefore, we develop techniques that can correct the biases introduced by noise. Correspondingly, a Newton-Raphson-type algorithm is proposed to obtain the estimator, leading to a corrected likelihood estimator. To further enhance computational efficiency, we introduce a corrected least squares estimator based on the idea of bias correction. These two estimation methods ensure both data security and the attainment of statistically valid estimators. Theoretical analysis of both estimators is carefully conducted, and statistical inference methods are discussed. The finite sample performances of different methods are demonstrated through extensive simulations and the analysis of a real dataset.
@@ 1. INTRODUCTION
A network refers to a set of nodes and their observed relationships (i.e., edges), and network data refer to the information collected from a network. Network data from different nodes are likely to be dependent on each other because different nodes in the networks are mutually connected with each other by edges. This is arguably the most important type of dependence induced by a network structure. For convenience, we refer to it as network dependence. To study network dependence, the spatial autoregressive (SAR, Ord 1975) model has been considered an extremely useful tool. The key idea of the SAR model is to assume that the response collected from each node is linearly related to its connected neighbors. This simple mechanism introduced a sophisticated and elegant network-dependency relationship, which leads to the frequent usage of the SAR model and its variants in real practice (Anselin et al., 2008; Chen et al., 2013; Zhu et al., 2017; Beenstock et al., 2019).
@@
In fact, the classical SAR models have inspired numerous follow-up studies. For example, Lee and Yu (2010) introduced the SAR panel model with individual effects and SAR disturbances. Yang and Lee (2017) and Zhu et al. (2020) developed a multivariate SAR model that models multivariate responses collected from each node. Various partially linear semiparametric spatial models were also developed to allow the SAR parameter to meaningfully vary across different nodes (Su, 2012; Malikova and Sun, 2017). The spatial dynamic panel data models were proposed to model time and network dependence simultaneously (Yu et al., 2008; Lee and Yu, 2014; Li, 2017). To compute the maximum likelihood estimator for various SAR models, the determinant of a large-scale matrix must be computed. This creates a significantly high level of computational complexity. To alleviate the computational cost, various smart computing algorithms were developed (Pace et al., 2000; Chen et al., 2013; Zhou et al., 2017). Particularly, the least squares method of Huang et al. (2019) and Zhu et al. (2020) has been demonstrated to be practically useful.
Notably, the above-mentioned studies assumed that the network data could be observed faithfully. This is indeed the case for many real applications where data privacy protection is not a serious issue. However, privacy protection has become of great concern in an increasing number of situations. Consider, for example, the “Wide-ranging Online Data for Epidemiologic Research” of the Centers for Disease Control and Prevention (CDC WONDER, https://wonder.cdc.gov/). This is a web-based tool for the dissemination of epidemiologic data collected by the National Vital Statistics System. Researchers can obtain details on the number of deaths attributed to different causes in a given region (e.g., county) and numerous auxiliary variables. Thus, each region can be treated as a node, and their spatial adjacency relationships become the edges. This leads to a network adjacency matrix, which is publicly known. We can also obtain accurate information related to particular regions, such as demographic data, economic development data, and healthcare coverage status. Ideally, various SAR models can be used to study the spatial dependence structure of lethal diseases. However, for many small counties, the number of deaths attributed to different causes together with auxiliary variables may be of significant privacy concern (Dinur and Nissim, 2003; Quick and Waller, 2018; Quick, 2021). From a broader perspective, in collaborations across different institutions, ensuring the protection of individual information necessitates privacy-protection measures for certain variables. Many similar scenarios exist, including medical research collaborations, transaction data analysis, and social media studies. To protect the privacy of individuals, appropriate measures have been developed and implemented to protect the raw data.
@@
These privacy protection measures include, but are not limited to, data swapping (Reiter, 2005), data imputation (Raghunathan et al., 2003; Reiter, 2005), posterior distribution sampling (Hu et al., 2018; Wilde et al., 2021), and noise addition (Dwork et al., 2006; Wilde et al., 2021; Ito et al., 2021). Among these measures, noise addition has gained arguably the greatest popularity. Its popularity may be due to two reasons. First, noise addition is a simple method and can be easily implemented in practice. Second, different levels of privacy protection can be provided by specifying different noise levels for the added random noise. In fact, the relationship between the added noise level and privacy-protection strength can be analytically studied using the theory of differential privacy (Dwork and Roth, 2014). However, extending this useful idea to privacy-protected spatial data analysis becomes a challenge. Therefore, we start with a classical SAR model with a faithfully observed response and publicly available explanatory variables. Subsequently, we consider protecting the raw response and some other explanatory variables with privacy-protection requirements by adding artificially generated random noise.
@@
In this study, we propose a privacy-protected SAR (PSAR) model to meet the privacy-protection requirements. Unlike previous SAR models, the unavailability of true values for both the response and some explanatory variables makes the classical quasi-maximum likelihood estimator (QMLE) inapplicable. To address this challenge, we first focus on deriving the likelihood function based on the true explanatory variables and observed responses. This enables the explicit expression of the log likelihood function and facilitates the derivation of first- and second-order derivatives. However, directly using observed explanatory variables may introduce a bias for the derivatives. Fortunately, the bias can be analytically expressed under suitable regularity conditions. Then, we compute the bias-corrected derivatives. Subsequently, a Newton-Raphson-type algorithm can be designed to obtain a corrected likelihood estimator (CLE). Nonetheless, computing the CLE becomes highly challenging for large-scale networks due to the computational complexity associated with high-dimensional matrix operations, such as determinant and inverse computations. To address this issue, we propose a corrected least squares estimator (CLS) for the PSAR model inspired by Huang et al. (2019), Zhu et al. (2020), and the concept of bias correction. Additionally, we propose a statistical inference method based on observed explanatory variables with noise added. In the theoretical analysis, we establish the identifiability, asymptotic normality, and numerical convergence properties for the estimators.
@@
The remainder of this paper is organized as follows. Section 2 describes the estimation methods in detail. Section 3 presents the theoretical properties from both statistical and numerical perspectives and provides a discussion of these methods. Section 4 presents the numerical results, including those of simulation studies and a real data example. Section 5 presents the concluding remarks. All technical details are relegated to the supplementary material.
@@ 2.1. Model and Notations
where $\rho_{0}\in\mathbb{R}$ is the network autocorrelation parameter measuring the network effect, and $\beta_{0}\in\mathbb{R}^{p}$ is the regression-coefficient vector reflecting the effect of nodal covariates. We define $S_{0}=I_{N}-\rho_{0}W$, where $I_{N}\in\mathbb{R}^{N\times N}$ is the identity matrix. We then have $\mathbb{Y}=S_{0}^{-1}(\mathbb{X}\beta_{0}+\mathbb{E})$. By omitting the constant term, the log-likelihood function for the parameter of interest $\theta=(\rho,\beta^{\top},\sigma^{2})^{\top}\in\mathbb{R}^{p+2}$ can be written as follows,
$\ell(\theta,\mathbb{Y},\mathbb{X})=\log|S|-N/2\log\sigma^{2}-1/2\sigma^{-2}\Big{(}S\mathbb{Y}-\mathbb{X}\beta\Big{)}^{\top}\Big{(}S\mathbb{Y}-\mathbb{X}\beta\Big{)}.$
Accordingly, a QMLE for $\theta$ can be obtained as $\hat{\theta}=\arg\max_{\theta}\ell(\theta,\mathbb{Y},\mathbb{X})$. Its asymptotic properties have been well studied. For example, refer to Lee (2004) and Anselin (2013).
However, in a privacy-protection scenario, we assume that the response $\mathbb{Y}$ and part of the explanatory covariates cannot be faithfully observed. Specifically, for the covariates, we assume that $\mathbb{X}=(\mathbb{X}_{1},\mathbb{X}_{2})$ with $\mathbb{X}_{1}\in\mathbb{R}^{N\times p_{1}}$, $\mathbb{X}_{2}\in\mathbb{R}^{N\times p_{2}}$, and $p_{1}+p_{2}=p$. Here, we define $\mathbb{X}_{1}$ to be the matrix collecting the faithfully observed covariates, and $\mathbb{X}_{2}^{*}=\mathbb{X}_{2}+\mathcal{E}_{x}\in\mathbb{R}^{N\times p_{2}}$ to be the matrix collecting explanatory variables with artificially created error matrix $\mathcal{E}_{x}=(\varepsilon_{x,ij})\in\mathbb{R}^{N\times p_{2}}$. Further, we assume that each element $\varepsilon_{x,ij}$ ($1\leq i\leq N,1\leq j\leq p_{2}$) in $\mathcal{E}_{x}$ is independently and identically distributed with mean 0 and predefined variance $\lambda_{x}^{2}$. Moreover, for the response, a data user can only observe $\mathbb{Y}^{*}=\mathbb{Y}+\mathcal{E}$ with an artificially created noise vector $\mathcal{E}=(\varepsilon_{1},\cdots,\varepsilon_{N})^{\top}\in\mathbb{R}^{N}$. Here, we assume that different $\varepsilon_{i}$s are independent and identically distributed with mean 0 and known variance $\lambda^{2}$. We assume here that data users cannot observe the true response variables $\mathbb{Y}$ and some explanatory variables $\mathbb{X}_{2}$. However, to better utilize the data, they are aware of the privacy-protection parameters $\lambda^{2}$ and $\lambda_{x}^{2}$. This assumption ensures the privacy of the data while simultaneously facilitating better estimation results for data users.
@@
where $\beta_{01}\in\mathbb{R}^{p_{1}}$ represents the coefficient for the truly observed covariates, $\beta_{02}\in\mathbb{R}^{p_{2}}$ is the coefficient corresponding to covariates with noise added, and we have $\beta_{0}=(\beta_{01}^{\top},\beta_{02}^{\top})^{\top}$. For convenience, we refer to this model in (2.2) as the privacy-protected spatial autoregressive (PSAR) model. Notably, model (2.2) only represents the relationship between the observable $\mathbb{Y}^{*}$ and $\mathbb{X}^{*}$. However, the challenge here is that $\mathbb{Y}^{*}$ is generated by true $\mathbb{X}$, which is not fully observable to data users.
@@ 2.2. Corrected Likelihood Estimator
By (2.3), the negative log-likelihood function for $\mathbb{Y}^{*}$ could be expressed as,
@@
where $\Omega=\sigma^{2}I_{N}+\lambda^{2}SS^{\top}$. Ideally, we can apply the classical Newton-Raphson algorithm to solve the optimization problem $\arg\min_{\theta}\mathcal{L}(\theta)$, and obtain estimates with desirable properties. Unfortunately, equation (2.4) is computationally infeasible in real practice because we can only observe $\mathbb{X}_{2}^{*}$. One solution is to directly substitute $\mathbb{X}_{2}$ with $\mathbb{X}_{2}^{*}$ in (2.4) for feasible computation, which yields $\mathcal{L}^{*}(\theta)=\mathcal{L}(\theta,\mathbb{Y}^{*},\mathbb{X}^{*})$. However, this approach will inevitably result in biased estimation. This is because the key statistical estimators derived from $\mathcal{L}^{*}(\theta)$, including the first and second derivatives, are all biased. We then analyze these derivatives to confirm this finding, and subsequently develop a bias-corrected Newton-Raphson-type algorithm to address this problem.
We first analyze the first derivatives of $\mathcal{L}(\theta)$ and $\mathcal{L}^{*}(\theta)$. To this end, we define ˙$\dot{\mathcal{L}}(\theta)=\Big{(}\dot{\mathcal{L}}^{\rho}(\theta),\big{\{}\dot{\mathcal{L}}^{\beta}(\theta)\big{\}}^{\top},\dot{\mathcal{L}}^{\sigma^{2}}(\theta)\Big{)}^{\top}\in\mathbb{R}^{p+2}$ to be the first-order derivative of $\mathcal{L}(\cdot)$ with respect to $\theta$, where ˙$\dot{\mathcal{L}}^{\rho}(\theta)$, ˙$\dot{\mathcal{L}}^{\beta}(\theta)$, and ˙$\dot{\mathcal{L}}^{\sigma^{2}}(\theta)$ are the first-order derivatives of $\mathcal{L}(\theta)$ with respect to $\rho$, $\beta$, and $\sigma^{2}$, respectively. Then, we can calculate that,
˙$\dot{\mathcal{L}}^{\rho}(\theta)$ $=$ $-\lambda^{2}\mbox{tr}\Big{(}\Omega^{-1}WS^{\top}\Big{)}+\lambda^{2}\widetilde{\mathbb{V}}^{\top}\Omega^{-1}WS^{\top}\Omega^{-1}\widetilde{\mathbb{V}}+\mbox{tr}\Big{(}WS^{-1}\Big{)}-\widetilde{\mathbb{V}}^{\top}\Omega^{-1}W\mathbb{Y}^{*},$
˙$\dot{\mathcal{L}}^{\beta}(\theta)$ $=$ $-\mathbb{X}^{\top}\Omega^{-1}\widetilde{\mathbb{V}},~{}\mbox{and}~{}~{}~{}\dot{\mathcal{L}}^{\sigma^{2}}(\theta)=\frac{1}{2}\mbox{tr}\Big{(}\Omega^{-1}\Big{)}-\frac{1}{2}\widetilde{\mathbb{V}}^{\top}\Omega^{-2}\widetilde{\mathbb{V}},$ (2.5)
where $\widetilde{\mathbb{V}}=S\mathbb{Y}^{*}-\mathbb{X}\beta$, and we should have $E\big{\{}\dot{\mathcal{L}}(\theta_{0})\big{\}}=\mathbf{0}_{p+2}$, where $\mathbf{0}_{p}$ is a $p$-dimensional zero vector. Subsequently, we define ˙$\dot{\mathcal{L}}^{*}(\theta)$ to be the first-order derivative of $\mathcal{L}^{*}(\theta)$, and we can observe $E\big{\{}\dot{\mathcal{L}}^{*}(\theta_{0})\big{\}}\neq 0$. To be more precise, let ˙$\dot{\mathcal{L}}^{*}(\theta)=\Big{(}\dot{\mathcal{L}}^{*\rho}(\theta),\dot{\mathcal{L}}^{*\beta}(\theta)^{\top},\dot{\mathcal{L}}^{*\sigma^{2}}(\theta)\Big{)}^{\top}\in\mathbb{R}^{p+2}$, where ˙$\dot{\mathcal{L}}^{*\beta}(\theta)=\big{(}\dot{\mathcal{L}}^{*\beta_{1}}(\theta)^{\top},\dot{\mathcal{L}}^{*\beta_{2}}(\theta)^{\top}\big{)}^{\top}$. It could be computed that,
@@
$E\big{\{}\dot{\mathcal{L}}^{*\sigma^{2}}(\theta_{0})\big{\}}$ $=$ $-\frac{1}{2}\lambda_{x}^{2}\beta_{02}^{\top}\beta_{02}\mbox{tr}(\Omega_{0}^{-2}).$ (2.8)
We define this difference for the first-order derivative caused by observed $\mathbb{X}^{*}$ as $\Delta\mathbf{S}_{\rm CL}(\theta_{0})=E\big{\{}\dot{\mathcal{L}}^{*}(\theta_{0})\big{\}}$.
Similarly, we could define the second order derivative for $\mathcal{L}(\theta)$ and $\mathcal{L}^{*}(\theta)$, which is denoted by ¨$\ddot{\mathcal{L}}(\theta)$ and ¨$\ddot{\mathcal{L}}^{*}(\theta)$, respectively. Define $\mathbb{W}_{S}=WS^{\top}+SW^{\top}$. Then we have,
¨$\ddot{\mathcal{L}}^{\rho\rho}(\theta)$ $=$ $-\lambda^{4}\mbox{tr}\Big{(}\Omega^{-1}\mathbb{W}_{S}\Omega^{-1}WS^{\top}\Big{)}+\lambda^{4}\widetilde{\mathbb{V}}^{\top}\Omega^{-1}\mathbb{W}_{S}\Omega^{-1}\mathbb{W}_{S}\Omega^{-1}\widetilde{\mathbb{V}}$
$-\lambda^{2}\widetilde{\mathbb{V}}^{\top}\Omega^{-1}WW^{\top}\Omega^{-1}\widetilde{\mathbb{V}}-2\lambda^{2}\widetilde{\mathbb{V}}^{\top}\Omega^{-1}\mathbb{W}_{S}\Omega^{-1}W\mathbb{Y}^{*}$
$+\lambda^{2}\mbox{tr}\Big{(}\Omega^{-1}WW^{\top}\Big{)}+\mbox{tr}\Big{(}WS^{-1}WS^{-1}\Big{)}+\mathbb{Y}^{*\top}W^{\top}\Omega^{-1}W\mathbb{Y}^{*},$
¨$\ddot{\mathcal{L}}^{\rho\beta}(\theta)$ $=$ $-\lambda^{2}\mathbb{X}^{\top}\Omega^{-1}\mathbb{W}_{S}\Omega^{-1}\widetilde{\mathbb{V}}+\mathbb{X}^{\top}\Omega^{-1}W\mathbb{Y}^{*},$ (2.9)
¨$\ddot{\mathcal{L}}^{\rho\sigma^{2}}(\theta)$ $=$ $\frac{\lambda^{2}}{2}\mbox{tr}\Big{(}\Omega^{-1}\mathbb{W}_{S}\Omega^{-1}\Big{)}+\mathbb{Y}^{*\top}W^{\top}\Omega^{-2}\widetilde{\mathbb{V}}-\lambda^{2}\widetilde{\mathbb{V}}^{\top}\Omega^{-1}\mathbb{W}_{S}\Omega^{-2}\widetilde{\mathbb{V}},$
¨$\ddot{\mathcal{L}}^{\beta\beta^{\top}}(\theta)$ $=$ $\mathbb{X}^{\top}\Omega^{-1}\mathbb{X},~{}~{}~{}\ddot{\mathcal{L}}^{\beta\sigma^{2}}(\theta)=\mathbb{X}^{\top}\Omega^{-2}\widetilde{\mathbb{V}},~{}~{}~{}\ddot{\mathcal{L}}^{\sigma^{2}\sigma^{2}}(\theta)=-\frac{1}{2}\mbox{tr}(\Omega^{-2})+\widetilde{\mathbb{V}}^{\top}\Omega^{-3}\widetilde{\mathbb{V}}.$
Similarly, ¨$\ddot{\mathcal{L}}^{*}(\theta)$ could be calculated with $\mathbb{X}$ replaced by $\mathbb{X}^{*}$. Then, define the difference between the expectations of second order derivatives for $\mathcal{L}^{*}(\theta_{0})$ and $\mathcal{L}(\theta_{0})$ as,
@@
It could be verified that $\Delta\mathbf{H}_{\rm CL}$ is symmetric and

### v2

@@ Abstract
Spatial autoregressive (SAR) models are important tools for studying network effects. However, with an increasing emphasis on data privacy, data providers often implement privacy protection measures that make classical SAR models inapplicable. In this study, we introduce a privacy-protected SAR model with noise-added response and covariates to meet privacy-protection requirements. However, in this scenario, the traditional quasi-maximum likelihood estimator becomes infeasible because the likelihood function cannot be directly formulated. To address this issue, we first consider an explicit expression for the likelihood function with only noise-added responses. Then, we develop techniques to correct the biases for derivatives introduced by noise. Correspondingly, a Newton-Raphson-type algorithm is proposed to obtain the estimator, leading to a corrected likelihood estimator. To further enhance computational efficiency, we introduce a corrected least squares estimator based on the idea of bias correction. These two estimation methods ensure both data security and the attainment of statistically valid estimators. Theoretical analysis of both estimators is carefully conducted, statistical inference methods and model extensions are discussed. The finite sample performances of different methods are demonstrated through extensive simulations and the analysis of a real dataset.
@@ 1. INTRODUCTION
A network refers to a set of nodes and their observed relationships (i.e., edges), and network data refers to the information collected from a network. Network data from different nodes are likely to be dependent on each other because different nodes in the networks are connected to each other by edges. This is arguably the most important type of dependence induced by a network structure. For convenience, we refer to it as network dependence. To study network dependence, the spatial autoregressive (SAR, Ord 1975) model has been considered an extremely useful tool. The key idea of the SAR model is to assume that the response collected from each node is linearly related to its connected neighbors. This simple mechanism introduces a sophisticated and elegant network-dependency relationship, which leads to the frequent use of the SAR model and its variants in real practice (Anselin et al., 2008; Chen et al., 2013; Zhu et al., 2017; Beenstock et al., 2019).
@@
In fact, the classical SAR models have inspired numerous follow-up studies. For example, Lee and Yu (2010) introduced the SAR panel model with individual effects and SAR disturbances. Yang and Lee (2017) and Zhu et al. (2020) developed a multivariate SAR model that models multivariate responses collected from each node. Various partially linear semiparametric spatial models were also developed to allow the SAR parameter to meaningfully vary across different nodes (Su, 2012; Malikova and Sun, 2017). The spatial dynamic panel data models were proposed to model time and network dependence simultaneously (Yu et al., 2008; Lee and Yu, 2014; Li, 2017). To compute the maximum likelihood estimator for various SAR models, the determinant of a large-scale matrix must be computed. This creates a significantly high level of computational complexity. To alleviate the computational cost, various efficient computing algorithms were developed (Pace et al., 2000; Chen et al., 2013; Zhou et al., 2017). In particular, the least squares method of Huang et al. (2019) and Zhu et al. (2020) has been demonstrated to be practically useful.
Notably, the aforementioned studies assume that network data can be observed accurately. This is true for many real-world applications where data privacy is not a major concern. However, in many situations, privacy protection has become increasingly important. Consider a data company that seeks collaboration with researchers for data analysis. While the company aims to achieve precise analysis results, it also wishes to protect the privacy of the data on its platform. For instance, a third-party payment platform may need researchers to analyze network effects among merchants on their platform. Network links, typically defined by merchants located in the same commercial district or in close proximity, can be directly observed and do not require privacy protection. However, the platform does not want to disclose specific business information of the merchants, such as turnover or average transaction value. In such cases, data sharing with researchers can only occur if there is a robust privacy protection mechanism in place. Similar scenarios are common in various fields, including medical research collaborations, transaction data analysis, and social media studies. To protect individual privacy, appropriate measures have been developed and implemented to safeguard raw data.
@@
These privacy protection measures include, but are not limited to, data swapping (Reiter, 2005), data imputation (Raghunathan et al., 2003; Reiter, 2005), posterior distribution sampling (Hu et al., 2018; Wilde et al., 2021), and noise addition (Dwork et al., 2006; Wilde et al., 2021; Ito et al., 2021). Among these measures, noise addition has arguably gained great popularity. Its popularity may be due to two reasons. First, noise addition is a simple method and can be easily implemented in practice. Second, different levels of privacy protection can be provided by specifying different noise levels for the added random noise. In fact, the relationship between the added noise level and privacy-protection strength can be analytically studied using the theory of differential privacy (Dwork and Roth, 2014). However, extending this useful idea to privacy-protected spatial data analysis becomes a challenge. When data companies share data, they aim to obtain accurate estimates while also protecting sensitive variables, whether the covariates $X$ or the response $Y$. This necessitates adding noise to both $X$ and $Y$. Therefore, we start with a classical SAR model and consider protecting the raw response and some explanatory variables that require privacy-protection by adding artificially generated random noise.
This issue is related to measurement errors in variables in traditional statistical research, where measurement errors in covariates have received considerable attention (Fuller, 1987; Stefanski and Carroll, 1987; Staudenmayer and Buonaccorsi, 2005; Carroll et al., 2006; Buonaccorsi, 2010; Hausman et al., 2021). Tools to deal with measurement error include, but are not limited to, simulation-extrapolation (Cook and Stefanski, 1994; Novick and Stefanski, 2002; Bertrand et al., 2017), instrumental variables (Schennach, 2007; Gustafson, 2007; Jiang and Ding, 2020; Luo et al., 2022), and the corrected score method (Stefanski and Carroll, 1987; Nakamura, 1990; Tsiatis and Ma, 2004; Wang et al., 2012). In the analysis of SAR models, scholars have also focused on addressing covariate measurement error. For example, Li et al. (2009) adopted structural model-based approaches to integrate out $X$ when the distribution of $X$’s measurement error is known. Luo et al. (2022) proposed the 3SLS method which uses instrumental variables to deal with measurement error in $X$. However, current literature on addressing measurement error in SAR models rarely considers scenarios where both $X$ and $Y$ contain noise, a common situation in privacy-preserving data analysis collaborations with data protection requirements. This might seem trivial in linear models because measurement errors in the continuous variable $Y$ do not affect the unbiasedness of the parameter estimates (Hausman, 2001). In SAR models, however, adding artificial noise to $Y$ inevitably leads to biased parameter estimates. Furthermore, in data collaborations, the sample size is typically large. However, existing literature on SAR with measurement errors seldom considers the computational complexity. Therefore, developing accurate and computationally efficient parameter estimation methods for SAR models in large-scale datasets for both noisy $X$ and $Y$ is of great importance in real practice.
@@
In this study, we propose a privacy-protected SAR (PSAR) model to meet the privacy-protection requirements. To obtain more accurate estimates, we assume that the data company has predetermined and disclosed the level of noise. Unlike previous SAR models, the unavailability of true values for both the response and some explanatory variables makes the classical quasi-maximum likelihood estimator (QMLE) inapplicable. To address this challenge, inspired by the corrected score method (Stefanski and Carroll, 1987; Nakamura, 1990), we first derive the likelihood function based on the true explanatory variables and observed responses. This approach allows for an explicit expression of the log-likelihood function and the derivation of its first- and second-order derivatives, which are, however, biased. Nonetheless, the bias can be analytically expressed under appropriate regularity conditions and subsequently corrected. Subsequently, a Newton-Raphson-type algorithm can be designed to obtain a corrected likelihood estimator (CLE). Nonetheless, computing the CLE becomes highly challenging for large-scale networks due to the computational complexity associated with high-dimensional matrix operations, such as determinant and inverse computations. To address this issue, we propose a corrected least squares estimator (CLS) for the PSAR model inspired by Huang et al. (2019), Zhu et al. (2020), and the concept of bias correction. Additionally, the statistical inference of the proposed method has been carefully discussed. In the theoretical analysis, we establish the identifiability, asymptotic normality, and numerical convergence properties for the estimators. Finally, we expand our analysis to encompass more complex scenarios. These include extending the classical SAR model to allow for multiple responses and multiple classes of node types in the networks, as well as incorporating perturbations to the network structure $W$ to enhance privacy protection.
@@
The remainder of this paper is organized as follows. Section 2 describes the estimation methods in detail. Section 3 presents the theoretical properties from both statistical and numerical perspectives and provides a discussion of generalization for these methods. Section 4 presents the numerical results, including those of simulation studies and a real data example. Section 5 presents the concluding remarks. All technical details are relegated to the supplementary material.
@@ 2.1. Model and Notations
where $\rho_{0}\in\mathbb{R}$ is the network autocorrelation parameter measuring the network effect, and $\beta_{0}\in\mathbb{R}^{p}$ is the regression-coefficient vector reflecting the effect of nodal covariates. We define $S_{0}=I_{N}-\rho_{0}W$, where $I_{N}\in\mathbb{R}^{N\times N}$ is the identity matrix. We then have $\mathbb{Y}=S_{0}^{-1}(\mathbb{X}\beta_{0}+\mathbb{E})$. By omitting the constant term, the log-likelihood function for the parameter of interest $\theta=(\rho,\beta^{\top},\sigma^{2})^{\top}\in\mathbb{R}^{p+2}$ can be obtained easily. Accordingly, a QMLE for $\theta$ can be obtained as $\hat{\theta}=\arg\max_{\theta}\ell(\theta,\mathbb{Y},\mathbb{X})$. Its asymptotic properties have been well studied. For example, refer to Lee (2004) and Anselin (2013).
However, in a privacy-protection scenario, we assume that the response $\mathbb{Y}$ and part of the explanatory covariates cannot be faithfully observed. Specifically, for the covariates, we assume that $\mathbb{X}=(\mathbb{X}_{1},\mathbb{X}_{2})$ with $\mathbb{X}_{1}\in\mathbb{R}^{N\times p_{1}}$, $\mathbb{X}_{2}\in\mathbb{R}^{N\times p_{2}}$, and $p_{1}+p_{2}=p$. Here, we define $\mathbb{X}_{1}$ to be the matrix collecting the faithfully observed covariates, and $\mathbb{X}_{2}^{*}=\mathbb{X}_{2}+\mathcal{E}_{x}\in\mathbb{R}^{N\times p_{2}}$ to be the matrix collecting explanatory variables with artificially created error matrix $\mathcal{E}_{x}=(\varepsilon_{x,ij})\in\mathbb{R}^{N\times p_{2}}$. Further, we assume that each element $\varepsilon_{x,ij}$ ($1\leq i\leq N,1\leq j\leq p_{2}$) in $\mathcal{E}_{x}$ is independently and identically distributed with mean 0 and predefined variance $\lambda_{x}^{2}$. Moreover, for the response, a data user can only observe $\mathbb{Y}^{*}=\mathbb{Y}+\mathcal{E}$ with an artificially created noise vector $\mathcal{E}=(\varepsilon_{1},\cdots,\varepsilon_{N})^{\top}\in\mathbb{R}^{N}$. Here, we assume that different $\varepsilon_{i}$s are independent and identically distributed with mean 0 and known variance $\lambda^{2}$. We assume here that data users cannot observe the true response variables $\mathbb{Y}$ and some explanatory variables $\mathbb{X}_{2}$. However, to better utilize the data, they are aware of the privacy-protection parameters $\lambda^{2}$ and $\lambda_{x}^{2}$. This assumption ensures the privacy of the data while simultaneously facilitating better estimation results for data users. Notably, if $\lambda^{2}$ is unknown, an identification issue arises for the parameter when $\rho=0$. However, since $\lambda^{2}$ is assumed to be known, this identification problem is eliminated.
@@
where $\beta_{01}\in\mathbb{R}^{p_{1}}$ represents the coefficient corresponding to the truly observed covariates, $\beta_{02}\in\mathbb{R}^{p_{2}}$ is the coefficient corresponding to those with noise added, and we have $\beta_{0}=(\beta_{01}^{\top},\beta_{02}^{\top})^{\top}$. For convenience, we refer to this model in (2.2) as the privacy-protected spatial autoregressive (PSAR) model. Notably, model (2.2) only represents the relationship between the observable $\mathbb{Y}^{*}$ and $\mathbb{X}^{*}$. However, the challenge here is that $\mathbb{Y}^{*}$ is generated by true $\mathbb{X}$, which is not fully observable to data users.
@@ 2.2. Corrected Likelihood Estimator
By (2.3), the negative log-likelihood function for $\theta$ based on $\mathbb{Y}^{*}$ and $\mathbb{X}$ is,
@@
where $\Omega=\sigma^{2}I_{N}+\lambda^{2}SS^{\top}$. Ideally, we can apply the classical Newton-Raphson algorithm to solve the optimization problem $\arg\min_{\theta}\mathcal{L}(\theta)$, and obtain estimates with desirable properties. Unfortunately, equation (2.4) is computationally infeasible in real practice because we can only observe $\mathbb{X}_{2}^{*}$. Direct substitution of $\mathbb{X}_{2}$ with $\mathbb{X}_{2}^{*}$ in (2.4) to obtain $\mathcal{L}^{*}(\theta)=\mathcal{L}(\theta,\mathbb{Y}^{*},\mathbb{X}^{*})$ for feasible computation will inevitably result in biased estimation. This is because the key statistical estimators derived from $\mathcal{L}^{*}(\theta)$, including the first- and second-order derivatives, are all biased.
We first analyze the difference between the first-order derivatives of $\mathcal{L}(\theta)$ and $\mathcal{L}^{*}(\theta)$. Define ˙$\dot{\mathcal{L}}(\theta)$ and ˙$\dot{\mathcal{L}}^{*}(\theta)$ to be the first-order derivative of $\mathcal{L}(\cdot)$ and $\mathcal{L}^{*}(\cdot)$ with respect to $\theta$, respectively. Then, we can calculate that $E\big{\{}\dot{\mathcal{L}}(\theta_{0})\big{\}}=\mathbf{0}_{p+2}$ and
@@
$E\big{\{}\dot{\mathcal{L}}^{*\sigma^{2}}(\theta_{0})\big{\}}$ $=$ $-\frac{1}{2}\lambda_{x}^{2}\beta_{02}^{\top}\beta_{02}\mbox{tr}(\Omega_{0}^{-2}),$ (2.7)
where ˙$\dot{\mathcal{L}}^{*\rho}(\theta)$, ˙$\dot{\mathcal{L}}^{*\beta}(\theta)$, and ˙$\dot{\mathcal{L}}^{*\sigma^{2}}(\theta)$ are the first-order derivatives of $\mathcal{L}(\theta)$ with respect to $\rho$, $\beta$, and $\sigma^{2}$, respectively. See Appendix A.1 for detailed verifications. We define this difference for the first-order derivative caused by $\mathbb{X}^{*}$ as $\Delta\mathbf{S}_{\rm CL}(\theta_{0})=E\big{\{}\dot{\mathcal{L}}^{*}(\theta_{0})\big{\}}-E\big{\{}\dot{\mathcal{L}}(\theta_{0})\big{\}}$.
Similarly, define the second-order derivative for $\mathcal{L}(\theta)$ and $\mathcal{L}^{*}(\theta)$ as ¨$\ddot{\mathcal{L}}(\theta)$ and ¨$\ddot{\mathcal{L}}^{*}(\theta)$, respectively. Define $\mathbb{W}_{S}=WS^{\top}+SW^{\top}$. In this way, the difference between the expectations of second-order derivatives for $\mathcal{L}^{*}(\theta_{0})$ and $\mathcal{L}(\theta_{0})$ could be calculated and defined as,
@@
It could be verified that $\Delta\mathbf{H}_{\rm CL}(\theta_{0})$ is symmetric and